# Dynamic Attention Analysis for Backdoor Detection in Text-to-Image Diffusion Models

### v2

@@ IV-F Ablation Study
Effect of start time step $t$ and time span $s$. Recall that in (7) and (17), $t$ and $s$ represent the start time step and time span, respectively. To evaluate their impact on DAA, we compute the F1 Score on the training set under different combinations of these parameters. As shown in Table IV, DAA-I achieves the best F1 Score when $t=3$ and $s=2$. DAA-S achieves the best F1 Score when $t=1$ and $s=5$. We observe that the anomalous dynamic feature of the backdoor attack is more effectively captured during the earlier stage of the diffusion process. We believe this is due to the relatively intense attention evolution during the early diffusion process. As shown in Fig. 5, the Lyapunov function derivative exhibits significant oscillations during the early 10 steps, implying that the evolution are more pronounced. The oscillation in early stages leads to more distinct dynamic features between benign and backdoor samples in the early stages.
@@
Effect of $\gamma_{L}$. Recall that $\gamma_{i},i\in[1,L]$ are the stability parameters for each node in the DAA-S dynamical system, with larger absolute values indicating greater stability. We perform an ablation experiment on $\gamma_{L}$ to evaluate its impact on detection performance. As shown in Table V, We observe that DAA-S exhibits similar F1 Scores and AUCs under different values of $\gamma_{L}$, demonstrating that the algorithm is insensitive to this parameter. Specifically, when $\gamma_{L}=-10$, the model achieves the best AUC value of 87.67%. Therefore, to obtain better feature extraction capability, we select $\gamma_{L}=-10$ as the optimal parameter.
@@
Effect of $c$. Recall that $c$ serves as the coupling coefficient in the dynamical system, balancing the influence between correlation features and node features within the system. Table VI presents the results of the ablation experiment. As shown, DAA-S achieves the best F1 Score of 81.88% when $c=5$, while achieving the best AUC of 87.67% when $c=1$. Finally, we select $c=1$ as the optimal parameter to achieve the best AUC performance.
@@
Effect of token selections. Here, we explore the impact of different token in calculating the Relative Evolve Rate (RER) in (5). Specifically, we consider two conditions: 1) using the $<$BOS$>$ token instead of the $<$EOS$>$ token, and 2) using the average Evolve Rate (ER) feature of all tokens. The average AUC for the five backdoor attack scenarios is shown in Table VII. As observed, the AUC for both conditions are around 65%, indicating that these features based on these tokens are not as effective as the features from $<$EOS$>$ token for backdoor detection. This result further highlights the unique role of the $<$EOS$>$ token in detecting backdoor samples.
DAA-S F1 Score (%) $\uparrow$ AUC (%) $\uparrow$ Avg F1 (%) $\uparrow$ Avg AUC (%) $\uparrow$
Rickrolling Villan EvilEdit IBA BadT2I Rickrolling Villan EvilEdit IBA BadT2I
[…]
F-norm 98.96 96.51 74.82 66.67 60.47 99.80 99.49 80.19 97.54 61.26 79.49 87.67
@@
Effect of similarity metric. In (10), we use the F-norm to represent the similarity between two attention maps. Here, we examine the impact of using 1-norm and SSIM [70] as similarity metrics on detection performance. As shown in Table VIII, the F-norm achieves the best performance, with an F1 score of 79.49% and an AUC of 87.67%. Therefore, we select the F-norm as the optimal similarity metric.
@@ IV-G Defense Robustness against Potential Adaptive Attack
In this section, we test the robustness of our methods against potential adaptive attacks. We assume that the attacker is aware of the backdoor detection mechanism of DAA, which leverages the anomalous attention of the $<$EOS$>$ token. The adaptive attack aims to avoid altering the attention of the $<$EOS$>$ token during backdoor injection. To achieve this, we propose a method based on EvilEdit [28]. Recall that EvilEdit injects backdoors into cross-attention via a closed-form solution:
@@
where $c^{tr}$ and $c^{ta}$ are the trigger embedding and backdoor target by CLIP text encoder [29], respectively. $W$ is the original cross-attention parameter, while $W^{*}$ is the edited parameter. $\lambda=1$ is the regularization hyper-parameter. $\mathbb{I}$ is the identity matrix. We perform the adaptive attack by removing the corresponding embedding of the $<$EOS$>$ token:
@@
where $L$ is the position of the $<$EOS$>$ token, $N=77$ is the token length. The same operation is applied to $c^{ta}$:
@@ V Conclusion
This paper introduces DAA, a novel approach for detecting backdoor samples in text-to-image diffusion models by leveraging dynamic attention analysis. Specifically, we find the dynamic anomaly of the $<$EOS$>$ token in backdoor samples and propose two effective detection methods based on this anomaly. Experimental results on five representative backdoor attack scenarios demonstrate the efficiency and effectiveness of DAA. Our work offers a new perspective on dynamic attention analysis to understand the behavior of backdoor samples. We believe our work opens up the floor for future investigation in both attacks and defenses, and contributes to build safe and reliable diffusion models.
@@
We provide the following supplementary materials in the Appendix, including the proof and additional details on our method.
@@
(PROPN)(,)(NOUN)(,)(VERB)(,)(NOUN)(NOUN)
@@
(NOUN)(,)(DET)(NOUN)(,)(NOUN)(,)(NOUN)
(,)(NOUN)(NOUN)(NOUN)

### v3

@@ IV-F Ablation Study
Effect of start time step $t$ and time span $s$. Recall that in Eq. (7) and Eq. (17), $t$ and $s$ represent the start time step and time span, respectively. To evaluate their impact on DAA, we compute the F1 Score on the training set under different combinations of these parameters. As shown in Table IV, DAA-I achieves the best F1 Score when $t=3$ and $s=2$. DAA-S achieves the best F1 Score when $t=1$ and $s=5$. We observe that the anomalous dynamic feature of the backdoor attack is more effectively captured during the earlier stage of the diffusion process. We believe this is due to the relatively intense attention evolution during the early diffusion process. As shown in Fig. 5, the Lyapunov function derivative exhibits significant oscillations during the early 10 steps, implying that the evolution are more pronounced. The oscillation in early stages leads to more distinct dynamic features between benign and backdoor samples in the early stages.
@@
Effect of $\gamma_{L}$. Recall that $\gamma_{i},i\in[1,L]$ are the stability parameters for each node in the DAA-S dynamical system, with larger absolute values indicating greater stability. We perform an ablation experiment on $\gamma_{L}$ to evaluate its impact on detection performance. As shown in Table V, We observe that DAA-S exhibits similar F1 Scores and AUCs under different values of $\gamma_{L}$, demonstrating that the algorithm is insensitive to this parameter. Here, we select $\gamma_{L}=-10$ as the optimal parameter to achieve the best balance between the F1 Score and AUC.
@@
Effect of $c$. Recall that $c$ serves as the coupling coefficient in the dynamical system, balancing the influence between correlation features and node features within the system. Table VI presents the results of the ablation experiment. As shown, DAA-S achieves the best F1 Score of 80.17% when $c=5$, while achieving the best AUC of 86.27% when $c=1$. Finally, we select $c=1$ as the optimal parameter to achieve the best AUC performance.
@@
Effect of token selections. Here, we explore the impact of different token in calculating the Relative Evolve Rate (RER) in Eq. (5). Specifically, we consider two conditions: 1) using the <<BOS>> token instead of the <<EOS>> token, and 2) using the average Evolve Rate (ER) feature of all tokens. The average AUC for the six backdoor attack scenarios is shown in Table VII. As observed, the AUC for both conditions are around 65%, indicating that these features based on these tokens are not as effective as the features from <<EOS>> token for backdoor detection. This result highlights the unique role of the <<EOS>> token in detecting backdoor samples.
@@
Effect of similarity metric. In Eq. (10), we use the F-norm to represent the similarity between two attention maps. Here, we examine the impact of using 1-norm and SSIM [78] as similarity metrics on detection performance. As shown in Table VIII, the F-norm achieves the best performance, with an F1 score of 79.27% and an AUC of 86.27%. Therefore, we select the F-norm as the optimal similarity metric.
@@ IV-G Defense Robustness against Potential Adaptive Attack
In this section, we test the robustness of our methods against potential adaptive attacks. We assume that the attacker is aware of the backdoor detection mechanism of DAA, which leverages the anomalous attention of the <<EOS>> token. The adaptive attack aims to avoid altering the attention of the <<EOS>> token during backdoor injection. To achieve this, we propose a method based on EvilEdit [28]. Recall that EvilEdit injects backdoors into cross-attention via a closed-form solution:
@@
where $c^{tr}$ and $c^{ta}$ are the trigger embedding and backdoor target by CLIP text encoder [29], respectively. $W$ is the original cross-attention parameter, while $W^{*}$ is the edited parameter. $\lambda=1$ is the regularization hyper-parameter. $\mathbb{I}$ is the identity matrix. We perform the adaptive attack by removing the corresponding embedding of the <<EOS>> token:
@@
where $L$ is the position of the <<EOS>> token, $N=77$ is the token length. The same operation is applied to $c^{ta}$:
@@ V Conclusion
This paper introduces DAA, a novel approach for detecting backdoor samples in text-to-image diffusion models by leveraging dynamic attention analysis. Specifically, we find the dynamic anomaly of the <<EOS>> token in backdoor samples and propose two effective detection methods based on this anomaly. Experimental results on six representative backdoor attack scenarios demonstrate the efficiency and effectiveness of DAA. Our work offers a new perspective on dynamic attention analysis to understand the behavior of backdoor samples. We believe our work opens up the floor for future investigation in both attacks and defenses, and contributes to build safe and reliable diffusion models.
@@
We provide the following supplementary materials in the Appendix, including the analysis and additional details on our method.
@@
(PROPN)(,)(NOUN)(,)(VERB)(,)$(NOUN)$
@@
(NOUN)(,)(DET)(NOUN)$(,)(NOUN)$
(,)(NOUN)$(NOUN)$
@@
PersonBA [v] motorbike bike
[v] bike motorbike
[v] tiger zebra
[v] zebra tiger
[v] dog cat
[v] cat dog